You have entered into a long forward contract on a dividend-paying stock some time ago, and this will expire in six months. It has a delivery price of $40 and the current stock price is $35. The stock provides a fixed dividend yield of 8% with semi-annual compounding. If the risk-free rate is 12% per annum with continuous compounding, what is the value of this long forward contract?

$6.72

-$4.02

$4.02

-$6.72
Answer #1

Value of long forward = spot rate / (1+dividend rate) ^ n - forward rate / e ^ r×n

= 35/ (1+0.08)^6/12 - 40/ e^ 0.12× 6/12

= 35 / (1.08)^0.5 - 40 / e^0.12 × 0.5

= 35 / 1.0392 - 40 / e^0.06

= 33.68 - 40 / 1.061

= 33.68 - 37.70

= -$4.02